OpenStudy (anonymous):

At a party, everyone shook hands with everybody else. There were 66 handshakes. How many people were at the party? Please Help, Thanks.

OpenStudy (neer2890):

for handshaking , if n people were there then total handshakes= n(n-1)

OpenStudy (anonymous):

Total number of people = (n-1)+(n-2)+(n-3)+(n-4)+...0 = (n-1)*(n-1+1)/2 = (n-1)*n/2 = 66 = n^2 -n = 132 =(n-12)(n+11) = 0; = n = 12 therefore 12 people

OpenStudy (anonymous):

Whoa! Thank you so much everyone! This study site is awesome! Thanks(:

OpenStudy (anonymous):

did you get it how to do it?

OpenStudy (anonymous):

Yes! I just need to look over the steps a few minutes and I'll get it(: I just needed like a base on how to get the answer(: Thanks.

OpenStudy (anonymous):

"Suppose there were 'n' people. Then each shook hands with (n-1) people. So total number of handshakes = n(n-1)/2 [Why divided by 2? The answer is whether person A shakes hand with person B or person B shakes hand with person A is immaterial as these two situations refer to a single handshake.] Given total no. of handshakes = 66 So n(n-1)/2 = 66 or n^2-n = 66*2 = 132 or n^2-n-132 = 0 or n^2-12n+11n-132 = 0 or n(n-12)+11(n-12) = 0 or (n-12)(n+11) = 0 Therefore either n = 12 or n = -11. But n being no. of people cannot be negative so n = 12 " " https://in.answers.yahoo.com/question/index?qid=1006060200769 "

OpenStudy (anonymous):

@no.name I thought we are not allowed to post answers from yahoo.

OpenStudy (anonymous):

Welcome to Openstudy

OpenStudy (anonymous):

@AntiNode I mentioned the link

OpenStudy (anonymous):

And if it is correct why not post

OpenStudy (anonymous):

@DivergentLover welcome.

OpenStudy (anonymous):

@No.name oh ok...

OpenStudy (anonymous):

Give me a reason , i checked the answer twice before posting

OpenStudy (anonymous):

Oh, Thank you too No.name :)

OpenStudy (anonymous):

Check this out :- (If you have time) Please http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

It was on the openstudyfeedback last time